A mixture of C3H8 and C2H2 has a mass of 2.9g. It is burned in excess O2 to form a mixture of water and carbon dioxide that contains 1.4 times as many moles of CO2 as of water.

Find the mass of C2H2 in the original mixture.

If possible, could I get a step by step solution for the problem?
Answer #1

The equation for the combustion of C3H8 and C2H2 can be written as follows.

2C3H8 + 4C2H2 + 19O2 ----> 14CO2 + 10H2O

i.e. The mole ratio C3H8:C2H2 = 2:4 = 1:2

i.e. x*44 + 2x*26 = 2.9

i.e. 44x + 52x = 2.9

i.e. 96x = 2.9

i.e. x = 2.9/96 = 0.03021 mol

Therefore, the mass of C2H2 in the mixture = 2*0.03021 mol * 26 g/mol = 1.571 g